# Mooncake: A KVCache-centric Disaggregated Architecture for LLM Serving

### v3

@@ 1.1 Motivation of Developing Mooncacke
Building on this idea, we found that the scheduling of KVCache is central to LLM serving scheduling. To improve overall throughput, there are typically two general approaches: 1) reuse KVCache as much as possible to reduce the required computation resources; and 2) maximize the number of tokens in each batch to improve the Model FLOPs Utilization (MFU). However, reusing KVCache from a remote location will prolong the TTFT, and a large batch size will lead to a larger TBT. Thus, the utilization of both these throughput oriented optimizations may lead to violations of latency related SLOs.
@@
In contrast, the decoding stage has different optimization goals and constraints. The aim is to aggregate as many tokens as possible in a decoding batch to improve MFU. However, this objective is restricted not only by the TBT SLO but also by the total size of aggregated KVCache that can be contained in the VRAM.
More importanly, existing research on LLM serving assumes sufficient resources and focuses on improving resource utilization. In contrast, the current GPU/accelerator supply is limited, and many MaaS providers face severe overload problems, especially during peak times. Scheduling in such scenarios presents unique challenges that existing works have not explored. For example, we need to predict future loads and reject certain requests early if there will be no available decoding slots after the prefill stage, to save wasted computation resources. However, a straightforward implementation of such an early reject policy surprisingly leads to fluctuations in the overloads. This has led us to aim at predicting the generation length of specific queries and making overall load predictions in the short-term future to implement a better rejection policy. It is also necessary to classify different request priorities to implement priority-based scheduling. In this paper, we summarize these problems as overload-oriented scheduling and present our preliminary study results.
@@ 1.2 Design and Results of Mooncacke
[…]
exampledefaultlinecolor=black,linewidth=3pt,leftmargin=1cm,rightmargin=1cm {mdframed}[style=exampledefault] To protect proprietary information and facilitate reproducibility, all the experimental results reported in this paper are based on replayed traces of real workloads, but using a dummy model that follows the same architecture as LLaMA2-70B. The trace includes only the timing of request arrivals, the number of input tokens, and the number of output tokens, the remmaped block hash, without any real user content. The trace is open-sourced at https://github.com/kvcache-ai/Mooncake later after following certain internal procedures.
@@ 2 Preliminary and Problem Definition
Modern large language models (LLMs) are based on the Transformer architecture, which utilizes attention mechanisms and multilayer perceptrons (MLP) to process input. Popular Transformer-based models, such as GPT [10] and LLaMA [11], employ a decoder-only structure. Each inference request is logically divided into two stages: the prefill stage and the decoding stage.
@@
3 Overview of Mooncake’s Disaggregated Archtecture
@@
Figure 3 illustrates the storage and transfer logic of the KVCache blocks. In CPU memory, KVCache is stored as paged blocks. Depending on the request patterns, it can use cache eviction algorithms such as LRU (Least Recently Used), LFU (Least Frequently Used), or algorithms based on request characteristics. The transfer of these KVCache blocks across CPUs and GPUs is handled by a separate (GPUDirect) RDMA-based component called Messenger. This archtecture also enables us to provide the context caching API to outside users for a higher reusement of KVCache.
@@
2) Incremental Prefill: The prefill node (group) completes the prefill stage using the prefix cache and stores the newly generated incremental KVCache back into CPU memory. If the number of uncached input tokens exceeds a certain threshold ($prefill\_chunk$), the prefill stage is split into multiple chunks and executed in a pipeline manner. This threshold is selected to fully utilize the corresponding GPU’s computational power and is typically larger than 1000 tokens. The reason for using chunked but still disaggregated prefill nodes is explained in §5.1.
@@
4 Sampled Real-wrold Request Trace
To facilitate further research on LLM serving, we sampled a subset of online request data from a 1-hour period to create an open-source request trace. To preserve caching relationships between requests, we prioritized collecting requests within the same session. The trace dataset comprises 23,608 entries, with fields including timestamp, input_length, output_length, and hash_ids. We included the remapped block hashes, which is particularly useful for analyzing and implementing KVCache reuse policies. To the best of our knowledge, this is the first open-source dataset that can be used for real-world reuse analysis.
@@ 4.1 Data Details
Hash ID: The hash_ids field describes prefix caching relationships. It is generated by hashing token blocks (with a block size of 512) into prefix hash values that include both the current and all preceding blocks (detailed in Figure 3). The resulting hash values are then mapped to globally unique IDs. Identical hash IDs indicate that a block of tokens, along with preceding tokens, are the same, thus allowing reuse within the corresponding KVCache. For example, in the provided samples, the first 12 hash IDs are identical, indicating they can share prefix caching for the first 12*512=6,144 tokens.
@@ 4.2 Statistical Features
We also conducted a simple cache policy analysis based on this trace, assuming a single global cache pool. Table 1 compares three cache strategies: LRU, LFU, and LengthAwareCache (similar to LFU but prioritizing eviction of cache blocks occurring later in requests) across different cache capacities. Increasing the cache capacity from 1,000 to 50,000 blocks boosts the cache hit ratio from 30% to 50%. Further capacity increases show minimal improvement. However, this should not be interpreted as an indication that larger caches are unnecessary, as the sample trace represents only a subset of real-world workloads. The required capacity should scale proportionally in actual scenarios. LRUCache performs best under this dataset’s patterns, likely due to the temporal proximity in request utilization.
@@ 5.1 Multi-node Prefill
To address this, Mooncake leverages the autoregressive property of decoder-only transformers and implements chunked pipeline parallelism (CPP) for long context prefill. We group every X nodes in the prefill cluster into a pipelined prefill node group. For each request, its input tokens are partitioned into chunks, each no longer than the $prefill\_chunk$. Different chunks of the same request can be processed simultaneously by different nodes, thus parallelizing the processing and reducing TTFT.
@@ 5.2 Layer-wise Prefill
Beyond computational power, the limited size of VRAM is also a precious resource, and we aim to minimize the VRAM occupation by states, primarily the KVCache. Theoretically, if the KVCache size of a request is $S$ and the processing time is $T$, its occupation cost is $S*T$. If a request is chunked and the processing of each chunk is inlined with other decoding requests in chunked prefill, $T$ will increase, leading to a larger occupation cost.
@@ 6 KVCache-centric Scheduling
[…]
15:         end if
16:    $\triangleright$ Cache-aware and -balancing prefill scheduling
[…]
@@
[…]
@@
[…]
@@ 6.1 Prefill Global Scheduling
Algorithm 1 details the mechanism for our cache-aware prefill scheduling. For every new request, its input tokens are divided into several blocks, and a hash key is computed for each block. This involves generating a hash key of tokens in a block concatenated with the hash key of the previous block (if available). The request’s block keys are then compared one by one against each prefill instance’s cache keys to identify the prefix match length ($prefix\_len$). Similar reuse logic is already implemented in vLLM, but the open-source version of vLLM only supports local KVCache caching.
With this matching information, Conductor estimates the corresponding execution time based on the request length and $prefix\_len$ (which varies by instance). It then adds the estimated waiting time for that request to get the TTFT on that instance. Finally, Conductor assigns the request to the instance with the shortest TTFT and updates the cache and queue times for that instance accordingly. If the SLO is not achievable, Conductor directly returns the HTTP 429 Too Many Requests response status code to the upper layers.
@@ 6.2 Cache Load Balancing
A straw-man solution to this KVCache scheduling problem could be collecting the global usages of each block, using a prediction model to forecast their future usages, and making scheduling decisions accordingly. However, unlike the estimation of prefill time, workloads are highly dynamic and change significantly over time. Especially for a MaaS provider experiencing rapid growth in its user base, it is impossible to accurately predict future usages. Thus, we propose a heuristic-based automated hot-spot migration scheme to enhance cache load balancing.
@@
As previously noted, requests may not always be directed to the prefill instance with the longest prefix cache length due to high instance load. In such cases, the conductor forwards the cache’s location and the request to an alternative instance if the estimated additional prefill time is shorter than the transfer time. This instance proactively retrieves the KVCache from the holder and stores it locally. More importantly, we prefer to compute the input tokens if the best remote prefix match length is no larger than the current local reusable prefix multiplied by a threshold [^1] This threshold is currently adjusted manually but can be adaptively adjusted by an algorithm in the future. Both strategies not only reduce the prefill time for requests but also facilitate the automatic replication of hot-spot caches, allowing for their broader distribution across multiple machines.
@@ 8.1 End-to-end Performance
Testbed During the experiments, the system was deployed on a high-performance computing node cluster to test performance. Each node in the cluster is configured as follows: 8 NVIDIA-A800-SXM4-80GB GPUs, each with 80GB HBM, connected by NVLINK; equipped with RDMA network cards that supporting up to 800 Gbps of interconnect bandwidth between nodes. Each node deploys either a prefill instance or a decoding instance according to the startup parameter.
@@
Metric In the experiments, we focus on the throughput performance of various systems under defined SLOs. We measure the TTFT and TBT across different RPS rates, where a higher RPS signifies improved throughput. To assess whether the majority of requests satisfy the SLOs, we use the 90th percentile (P90) values of TTFT and TBT as the ultimate metrics. As mentioned in §2, the thresholds for TTFT and TBT are set by multiplying the lowest observed RPS values by factors of 10 and 5, respectively. Exceeding these thresholds indicates a failure to meet the SLOs and the corresponding consumed resources are considered as wasted. For ease of comparison, we normalize all TTFT and TBT values against these upper limits, establishing a baseline of 1.0.
@@ 8.2 Performance in Overload Scenarios
In this section, we evaluate performance under overload scenarios, focusing on the maximum number of requests the system can handle, as discussed in §7. The baseline strategy, which rejects requests based on load before both stages start, leads to resource wastage by rejecting requests already processed in the prefill stage. In contrast, we propose the Early Rejection and Early Rejection based on Prediction strategies, detailed in §7.2 and §7.4, respectivly. These strategies take the system’s load into comprehensive consideration, and hence reducing unnecessary request rejections.
@@ 9 Related Work
Prefix caching is also widely adopted to enable the reuse of KV caches across multiple requests, reducing computational overhead in LLM inference systems [29, 13]. Prompt Cache [33] precomputes and stores frequently used text KV caches on inference servers, facilitating their reuse and significantly reducing inference latency. SGLang [34] leverages RadixAttention, which uses a least recently used (LRU) cache within a radix tree structure to efficiently enable automatic sharing across various reuse patterns.
Among these approaches, AttentionStore [35], a concurrent work with us, proposes a hierarchical KV caching system that utilizes cost-effective memory and storage media to accommodate KV caches for all requests. The architecture of Mooncake shares many design choices with AttentionStore. However, in long-context inference, the KV cache becomes extremely large, requiring high capacity and efficient data transfer along with KVCache-centric global scheduling. Additionally, Mooncake is not a standalone cache service, it incorporates both a memory-efficient cache storage mechanism and a cache-aware scheduling strategy, further improving prefix caching efficiency.
Furthermore, recent research [36] has started exploring the scheduling of prompts, which is essentially KVCache-centric scheduling. We corroborate many results in this area, although the real reusability in our online traces is much smaller than results reproduced by open-source benchmarks. Theoretically, up to only 50% of the KVCache can be reused in our current workloads, even if we assume both the capacity of storage and the TTFT SLO are infinite. However, this reusability highly depends on the application scenario and can be as large as 90% for certain scenarios, such as our chat-to-paper service https://papers.cool/. We also emphasize the need for overload-oriented scheduling subject to SLOs, rather than merely throughput-oriented scheduling.

### v4

@@ 1.1 Motivation of Developing Mooncacke
Building on this idea, we found that the scheduling of KVCache is central to LLM serving scheduling. To improve overall throughput, there are typically two general approaches: 1) reuse KVCache as much as possible to reduce the required computation resources; and 2) maximize the number of tokens in each batch to improve the Model FLOPs Utilization (MFU). However, reusing KVCache from a remote location will prolong the TTFT, and a large batch size will lead to a larger TBT. Thus, the utilization of both these throughput-oriented optimizations may lead to violations of latency-related SLOs.
@@
In contrast, the decoding stage has different optimization goals and constraints. The aim is to aggregate as many tokens as possible in a decoding batch to improve MFU. However, this objective is restricted not only by the TBT SLO but also by the total size of the aggregated KVCache that can be contained in the VRAM.
More importantly, existing research on LLM serving assumes sufficient resources and focuses on improving resource utilization. In contrast, the current GPU/accelerator supply is limited, and many MaaS providers face severe overload problems, especially during peak times. Scheduling in such scenarios presents unique challenges that existing works have not explored. For example, we need to predict future loads and reject certain requests early if there will be no available decoding slots after the prefill stage, to save wasted computation resources. However, a straightforward implementation of such an early reject policy surprisingly leads to fluctuations in the overloads. This has led us to aim at predicting the generation length of specific queries and making overall load predictions in the short-term future to implement a better rejection policy. It is also necessary to classify different request priorities to implement priority-based scheduling. In this paper, we summarize these problems as overload-oriented scheduling and present our preliminary study results.
@@ 1.2 Design and Results of Mooncacke
To protect proprietary information and facilitate reproducibility, all the experimental results reported in this paper are based on replayed traces of real workloads, but using a dummy model that follows the same architecture as LLaMA2-70B. The trace includes only the timing of request arrivals, the number of input tokens, and the number of output tokens, the remaped block hash, without any real user content. The trace is open-sourced at https://github.com/kvcache-ai/Mooncake.
@@ 2 Preliminary and Problem Definition
Modern large language models (LLMs) are based on the Transformer architecture, which utilizes attention mechanisms and multilayer perceptrons (MLPs) to process input. Popular Transformer-based models, such as GPT [10] and LLaMA [11], employ a decoder-only structure. Each inference request is logically divided into two stages: the prefill stage and the decoding stage.
@@
3 Overview of Mooncake’s Disaggregated Architecture
@@
Figure 3 illustrates the storage and transfer logic of the KVCache blocks. In CPU memory, KVCache is stored as paged blocks. Depending on the request patterns, it can use cache eviction algorithms such as LRU (Least Recently Used), LFU (Least Frequently Used), or algorithms based on request characteristics. The transfer of these KVCache blocks across CPUs and GPUs is handled by a separate (GPUDirect) RDMA-based component called Messenger. This architecture also enables us to provide the context caching API to outside users for a higher reuse of KVCache.
@@
2) Incremental Prefill: The prefill node (group) completes the prefill stage using the prefix cache and stores the newly generated incremental KVCache back into CPU memory. If the number of uncached input tokens exceeds a certain threshold (prefill_chunkprefill\_chunk), the prefill stage is split into multiple chunks and executed in a pipeline manner. This threshold is selected to fully utilize the corresponding GPU’s computational power and is typically larger than 1000 tokens. The reason for using chunked but still disaggregated prefill nodes is explained in §5.1.
@@
4 Sampled Real-world Request Trace
To facilitate further research on LLM serving, we sampled a subset of online request data from a 1-hour period to create an open-source request trace. To preserve caching relationships between requests, we prioritized collecting requests within the same session. The trace dataset comprises 23,608 entries, with fields including timestamp, input_length, output_length, and hash_ids. We included the remapped block hashes, which are particularly useful for analyzing and implementing KVCache reuse policies. To the best of our knowledge, this is the first open-source dataset that can be used for real-world reuse analysis.
@@ 4.1 Data Details
Hash ID: The hash_ids field describes prefix caching relationships. It is generated by hashing token blocks (with a block size of 512) into prefix hash values that include both the current and all preceding blocks (detailed in Figure 3). The resulting hash values are then mapped to globally unique IDs. Identical hash IDs indicate that a block of tokens, along with preceding tokens, is the same, thus allowing reuse within the corresponding KVCache. For example, in the provided samples, the first 12 hash IDs are identical, indicating they can share prefix caching for the first 12*512=6,144 tokens.
@@ 4.2 Statistical Features
We also conducted a simple cache policy analysis based on this trace, assuming a single global cache pool. Table 1 compares three cache strategies: LRU, LFU, and LengthAwareCache (similar to LFU but prioritizing cache blocks occurring later in requests) across different cache capacities. Increasing the cache capacity from 1,000 to 50,000 blocks boosts the cache hit ratio from 30% to 50%. Further capacity increases show minimal improvement. However, this should not be interpreted as an indication that larger caches are unnecessary, as the sample trace represents only a subset of real-world workloads. The required capacity should scale proportionally in actual scenarios. LRUCache performs best under this dataset’s patterns, likely due to the temporal proximity in request utilization.
@@ 5.1 Multi-node Prefill
To address this, Mooncake leverages the autoregressive property of decoder-only transformers and implements chunked pipeline parallelism (CPP) for long context prefill. We group every X nodes in the prefill cluster into a pipelined prefill node group. For each request, its input tokens are partitioned into chunks, each no longer than the prefill_chunkprefill\_chunk. Different chunks of the same request can be processed simultaneously by different nodes, thus parallelizing the processing and reducing TTFT.
@@ 5.2 Layer-wise Prefill
Beyond computational power, the limited size of VRAM is also a precious resource, and we aim to minimize the VRAM occupation by states, primarily the KVCache. Theoretically, if the KVCache size of a request is $S$ and the processing time is $T$, its occupation cost is STS*T. If a request is chunked and the processing of each chunk is inlined with other decoding requests in chunked prefill, $T$ will increase, leading to a larger occupation cost.
@@ 6 KVCache-centric Scheduling
7:for $\mathit{instance}\in P$ do
8:  $\mathit{prefix\_len}\leftarrow\mathit{instance.prefix\_len}$
9:  $\mathit{T_{queue}}\leftarrow\text{EstimatePrefillQueueTime}(\mathit{instance})$
10:  if $\frac{\mathit{best\_prefix\_len}}{\mathit{prefix\_len}}<{\bf kvcache\_balancing\_threshold}$ then $\triangleright$ Cache-aware prefill scheduling
11:    $\mathit{T_{prefill}}\leftarrow\text{EstimatePrefillExecutionTime}(\text{len}(R.\mathit{prompt\_tokens}),\mathit{prefix\_len})$
12:    if $\mathit{TTFT}>\mathit{T_{queue}}+\mathit{T_{prefill}}$ then
13:     $\mathit{TTFT}\leftarrow\mathit{T_{queue}}+\mathit{T_{prefill}}$
14:     $p\leftarrow\mathit{instance}$
15:    end if
@@
17:    $\mathit{transfer\_len}\leftarrow\mathit{best\_prefix\_len}-\mathit{prefix\_len}$
18:    $\mathit{T_{transfer}}\leftarrow\text{EstimateKVCacheTransferTime}(\mathit{instance},\mathit{best\_matched\_instance},\mathit{transfer\_len})$
19:    $\mathit{T_{prefill}}\leftarrow\text{EstimatePrefillExecutionTime}(\text{len}(R.\mathit{prompt\_tokens}),\mathit{best\_prefix\_len})$
20:    if $\mathit{TTFT}>\mathit{T_{transfer}}+\mathit{T_{queue}}+\mathit{T_{prefill}}$ then
21:     $\mathit{TTFT}\leftarrow\mathit{T_{transfer}}+\mathit{T_{queue}}+\mathit{T_{prefill}}$
22:     $p\leftarrow\mathit{instance}$
@@
27:if $\mathit{TTFT}>\mathit{TTFT\_SLO}$ or $\mathit{TBT}>\mathit{TBT\_SLO}$ then
@@
30:if $\frac{\mathit{best\_prefix\_len}}{p.\mathit{prefix\_len}}>{\bf kvcache\_balancing\_threshold}$ then
31:  $\text{TransferKVCache}(\mathit{best\_matched\_instance},p)$ $\triangleright$ KVCache hot-spot migration
@@ 6.1 Prefill Global Scheduling
Algorithm 1 details the mechanism for our cache-aware prefill scheduling. For every new request, its input tokens are divided into several blocks, and a hash key is computed for each block. This involves generating a hash key of tokens in a block concatenated with the hash key of the previous block (if available). The request’s block keys are then compared one by one against each prefill instance’s cache keys to identify the prefix match length (prefix_lenprefix\_len). Similar reuse logic is already implemented in vLLM, but the open-source version of vLLM only supports local KVCache caching.
With this matching information, Conductor estimates the corresponding execution time based on the request length and prefix_lenprefix\_len (which varies by instance). It then adds the estimated waiting time for that request to get the TTFT on that instance. Finally, Conductor assigns the request to the instance with the shortest TTFT and updates the cache and queue times for that instance accordingly. If the SLO is not achievable, Conductor directly returns the HTTP 429 Too Many Requests response status code to the upper layers.
@@ 6.2 Cache Load Balancing
A straw-man solution to this KVCache scheduling problem could be collecting the global usages of each block, using a prediction model to forecast their future usages, and making scheduling decisions accordingly. However, unlike the estimation of prefill time, workloads are highly dynamic and change significantly over time. Especially for a MaaS provider experiencing rapid growth in its user base, it is impossible to accurately predict future usage. Thus, we propose a heuristic-based automated hot-spot migration scheme to enhance cache load balancing.
@@
As previously noted, requests may not always be directed to the prefill instance with the longest prefix cache length due to high instance load. In such cases, the conductor forwards the cache’s location and the request to an alternative instance if the estimated additional prefill time is shorter than the transfer time. This instance proactively retrieves the KVCache from the holder and stores it locally. More importantly, we prefer to compute the input tokens if the best remote prefix match length is no larger than the current local reusable prefix multiplied by a threshold [^1] This threshold is currently adjusted manually, but can be adaptively adjusted by an algorithm in the future. Both strategies not only reduce the prefill time for requests but also facilitate the automatic replication of hot-spot caches, allowing for their broader distribution across multiple machines.
@@ 8.1 End-to-end Performance
Testbed During the experiments, the system was deployed on a high-performance computing node cluster to test performance. Each node in the cluster is configured as follows: 8 NVIDIA-A800-SXM4-80GB GPUs, each with 80GB HBM, connected by NVLINK; equipped with RDMA network cards that support up to 800 Gbps of interconnect bandwidth between nodes. Each node deploys either a prefill instance or a decoding instance according to the startup parameter.
@@
Metric In the experiments, we focus on the throughput performance of various systems under defined SLOs. We measure the TTFT and TBT across different RPS rates, where a higher RPS signifies improved throughput. To assess whether the majority of requests satisfy the SLOs, we use the 90th percentile (P90) values of TTFT and TBT as the ultimate metrics. As mentioned in §2, the thresholds for TTFT and TBT are set by multiplying the lowest observed RPS values by factors of 10 and 5, respectively. Exceeding these thresholds indicates a failure to meet the SLOs, and the corresponding consumed resources are considered as wasted. For ease of comparison, we normalize all TTFT and TBT values against these upper limits, establishing a baseline of 1.0.
@@ 8.2 Performance in Overload Scenarios
In this section, we evaluate performance under overload scenarios, focusing on the maximum number of requests the system can handle, as discussed in §7. The baseline strategy, which rejects requests based on load before both stages start, leads to resource wastage by rejecting requests already processed in the prefill stage. In contrast, we propose the Early Rejection and Early Rejection based on Prediction strategies, detailed in §7.2 and §7.4, respectively. These strategies take the system’s load into comprehensive consideration, and hence reduce unnecessary request rejections.
@@ 9 Related Work
Prefix caching is also widely adopted to enable the reuse of KVCache across multiple requests, reducing computational overhead in LLM inference systems [29, 13]. Prompt Cache [33] precomputes and stores frequently used text KVCache on inference servers, facilitating their reuse and significantly reducing inference latency. SGLang [34] leverages RadixAttention, which uses a least recently used (LRU) cache within a radix tree structure to efficiently enable automatic sharing across various reuse patterns.
Among these approaches, AttentionStore [35], a concurrent work with us, proposes a hierarchical KVCache system that utilizes cost-effective memory and storage media to accommodate KVCache for all requests. The architecture of Mooncake shares many design choices with AttentionStore. However, in long-context inference, the KVCache becomes extremely large, requiring high capacity and efficient data transfer along with KVCache-centric global scheduling. Additionally, Mooncake is not a standalone cache service; it incorporates both a memory-efficient cache storage mechanism and a cache-aware scheduling strategy, further improving prefix caching efficiency.
Furthermore, recent research [36] has started exploring the scheduling of prompts, which is essentially KVCache-centric scheduling. We corroborate many results in this area, although the real reusability in our online traces is much smaller than the results reproduced by open-source benchmarks. Theoretically, up to only 50% of the KVCache can be reused in our current workloads, even if we assume both the capacity of storage and the TTFT SLO are infinite. However, this reusability highly depends on the application scenario and can be as large as 90% for certain scenarios, such as our chat-to-paper service https://papers.cool/. We also emphasize the need for overload-oriented scheduling subject to SLOs, rather than merely throughput-oriented scheduling.